# On the Constant Depth Implementation of Pauli Exponentials

### v5

@@ Abstract
We decompose for the first time, under the very restrictive linear nearest-neighbour connectivity, $Z\otimes Z\ldots\otimes Z$ exponentials of arbitrary length into circuits of constant depth using $\mathcal{O}(n)$ ancillae and two-body XX and ZZ interactions. Consequently, a similar method works for arbitrary Pauli exponentials. We prove the correctness of our approach, after introducing novel rewrite rules for circuits which benefit from qubit recycling. The decomposition has a wide variety of applications ranging from the efficient implementation of fault-tolerant lattice surgery computations, to expressing arbitrary stabilizer circuits via two-body interactions only, parallel decoding of quantum error-correcting computations and to reducing the depth of NISQ computations, such as VQE.
@@ I Introduction
We introduce a novel decomposition of Pauli exponentials, which uses a linear number of ancillas (effectively doubling the number of qubits) and only two-body XX and ZZ interactions. Two-body interactions are the easiest to engineer and implement, and are native to Majorana computers [2], ion traps [10], silicon spin qubits [11], as well as to fault-tolerant and error-correction circuits. In other words, we present a very efficient way of emulating the all-to-all connectivity (on linear nearest-neighbour architectures) required for executing exponentiated Pauli strings.
@@
Figure 1: An $n$-qubit Pauli exponential is decomposed into a constant depth sequence of XX and ZZ interactions and $\mathcal{O}(n)$ ancillae. We use the shorthand notation $e^{i\theta P}=P(\theta)$, and consider that $P$ is formed exclusively of $Z$ terms: $P=Z\otimes Z\ldots\otimes Z$. A general string can always be decomposed into single-qubit Clifford gates conjugating the $Z$-terms of $P$ [9]. In this figure, we assume that $\theta$ is arbitrary, such that our construction is extending $P$ with an additional $Z$ term (marked green) and an ancilla initialized in |$\ket{\theta}=Rz(\theta)\ket{+}$ [4].
@@ I.1 Novelty
Our decomposition is applicable to both NISQ and error-corrected machines. In contrast, a result concerning long range CNOTs has been proposed by [19]. Their adaptive circuit construction uses the same number of ancillae as our decomposition. Nevertheless, due to the exclusive focus on CNOT gates, their construction is only suitable for NISQ computers. Another constant-depth implementation of long range CNOTs using pairwise measurements is presented in [16], and is motivated by the fact that efficient decompositions of Pauli exponentials were not known at the time. Our work bridges the gap that motivated [16].
@@ II Methods
The lhs. of Fig. 1 represents the unitary obtained by the exponentiation. The rhs. is the decomposition which includes ancillae for enabling the XX interactions, as well as the |$\ket{\theta}$ ancilla initialized for performing a teleportation-based gate.
Figure 2: The ROT rewrite rule implements the exponential of the Pauli $P$ by introducing an ancilla initialized in |$\ket{\theta}=Rz(\theta)\ket{+}$, extending the Pauli string with a $Z$-term for the ancilla, and measuring the ancilla in the X-basis[4].
@@ II.1 Measurement-based implementation
When implementing the Pauli exponentials on fault-tolerant machines, the $\theta$ rotation is first decomposed into a sequence of rotation gates supported by the QEC. This increases the depth proportionally to the length of the gate sequence. Moreover, the depth will additionally increase by a constant factor, as each rotation is probabilistic and might require corrections [4] (e.g. if the X-measurement results in the -1 eigenvalue, then a corrective rotation of $2\theta$ is needed). Corrections can be tracked in a Pauli frame, when the angles correspond to Pauli gates.
Moreover, we assume that the two-body interactions are also measurement-based. The latter will introduce corrective terms (e.g. [3]), but the corrections can be tracked in a Pauli frame, as the underlying circuit will consist solely of CNOT gates and ancillae initialised or measured in a discrete set of rotated basis [20, 21].
@@ II.2 Decomposing the Pauli exponentials
We start by decomposing an arbitrary Pauli string $P$ into a string $C=C_{0}\otimes\ldots\otimes C_{n}$ of single qubit Clifford gates $C_{i}$, and a string $P_{z}=Z\otimes\ldots\otimes Z$, such that $P=CP_{z}C^{\dagger}$. As an example, single qubit gates can be efficiently implemented in practice [22, 23] in surface codes. We continue by decomposing $P_{z}$ via the PG rule [9] (Fig. 3). CNOTs are expressed in lattice surgery style[24] decompositions (Fig. 4). After using PG, we choose the first LS rewrite for the lhs. CNOT, and the second rewrite for the rhs. CNOT.
@@
We replace a sequence of measurement and reset operations in the same basis with an equivalent measurement-based implementation by using the MR rule (Fig. 5). This is, to the best of our knowledge, a novel quantum circuit rewrite rule. The rule is the inverse of the copy rule from the ZX-calculus [25].
@@ III.1 Lattice Surgery
Lattice surgery is the de facto standard way of implementing computations with QLDPC codes [6], such as the surface code. Therein, the logical qubits of the circuits are arranged on a 2D layout of patches (e.g. Fig. 7), and ancillary space between the patches supports long range multi-body measurements. The latter are expressed as generalized Pauli exponentials. The structure of the layout plays a significant role in the trade-off between a computation’s speed and the required amount of physical hardware [28, 4]. Although our decomposition uses ancillae, these are readily available in most of the lattice surgery layouts considered for compiling fault-tolerant computations.
Figure 7: Multi-body $Z\otimes\ldots\otimes Z$ interaction implemented by lattice surgery. The patch layout is adapted from  [28]: pink patches (Q) are holding logical qubits, and gray patches (A) are ancilla for XX and ZZ interactions. The blue boundaries are the Z operators of the logical A patches. a) according to Fig. 1, we initialize two ancilla in |$\ket{0}$ and perform XX two-body interactions (orange) with the Q patches; c) afterwards we perform pairwise ZZ two-body interactions (blue) between pairs of Q and A patches. Finally, the interactions from a) are repeated. The interaction schedule from a) and b) is highly regular and can be further optimized in terms of hardware overheads.
The general compilation strategy for lattice surgery is to first decompose arbitrary algorithms into generalized Pauli exponentials [4], and then to implement these exponentials via multi-body measurements. This approach has multiple disadvantages.
@@ III.2 Parallel Decoding
Our decomposition enables the parallelisation of decoding by setting an upper bound on the maximum (equivalent) decodable distance that a QEC decoder will need to handle within one QEC cycle. In the following, we introduce the concept of equivalent decodable distance and restrict the discussion, without loss of generality, to rotated surface codes operated by lattice surgery.
@@
At the same time, as seen in Fig. 7, all blue patches can be decoded in parallel. Our parallelisation method is reminiscent of the proposal from [31] (fixing the number of physical qubits per decoder) and the constructions from [32, 33] (decoders which communicate messages). For the latter, the decoders will communicate the observed correlated logical errors [17] after the long-range interaction is split via lattice surgery.
@@ III.3 Constant Depth Fault-Tolerant Protocols
From the perspective of stabilizer QECC codes, our decomposition allows any set of QECC stabilizers to be implemented in constant depth using pairwise XX and ZZ measurements. QECC code conversion schemes [34] can also be implemented locally using pairwise interactions. Additionally, after carefully choosing the pairwise interaction schedules, our method might lead to results similar to the ones from [35]. Our decomposition involves only two-body measurements and may find relevance to subsystem codes such as the Bacon-Shor code [36] or its Floquet version [35].
@@ IV Generalization
So far, we considered constant depth decompositions of Pauli strings of the form $Z\otimes...\otimes Z$. Nevertheless, there might be qubits which are not involved in the operation, and the corresponding Pauli string will include an $I$ term for those qubits.
@@ IV.1 Qudits
An alternative to using ququarts is to use one additional qubit $q$ placed next to each ancilla $a$ from Fig. 1. These $q$ qubits will be initialized in |$\ket{0}$, and the $I$ operators from the Pauli string would be replaced by $SWAP(a,q)$, a $Z$ on $a$, followed by $SWAP(a,q)$. This construction would not be nearest-neighbour, but next-nearest neighbour.
@@ IV.2 The qubit layout
In a linear nearest neighbour layout of qubits, the logical qubits are arranged in a line similar to $\begin{smallmatrix}Q&0&Q&0\ldots&Q\end{smallmatrix}$, where 00 are ancillae initialized in |$\ket{0}$. In this scenario, pairwise nearest-neighbour interactions are not possible without resorting to ququarts, as discussed in the previous section. Consequently, the long range construction from [19] uses nearest-neighbour pairwise interactions, but it uses the less general layout $\begin{smallmatrix}Q&0&0&0\ldots&Q\end{smallmatrix}$, where all the qubits between the first and last one are initialized in |$\ket{0}$.
A bi-linear nearest neighbour layout of qubits of the form $\begin{smallmatrix}Q&0&Q&0\ldots&Q\\ r&r&r&r\ldots&r\end{smallmatrix}$ allows for jumping over qubits operated by $I$ and maintaining the nearest-neighbour decomposition – the routing qubits $r$ can be used first to prepare Bell pairs using the constant depth recipe from [16]. The layout from Fig. 7 has the sane properties the bi-linear nearest neighbour layout has.

### v6

@@ Abstract
We decompose, under the very restrictive linear nearest-neighbour connectivity, $Z^{\otimes n}$ exponentials of arbitrary length into circuits of constant depth using $\mathcal{O}(n)$ ancillae and two-body XX and ZZ interactions. Consequently, a similar method works for arbitrary Pauli exponentials. We prove the correctness of our approach, after introducing novel rewrite rules for circuits which benefit from qubit recycling. The decomposition has a wide variety of applications ranging from the efficient implementation of practical fault-tolerant lattice surgery computations, to expressing arbitrary stabilizer circuits via two-body interactions only and parallel decoding of quantum error-correcting computations.
@@ I Introduction
We introduce a novel decomposition of Pauli exponentials, which uses a linear number of ancillas (effectively doubling the number of qubits) and only two-body XX and ZZ interactions. Depending on where the Pauli exponentials are executed, the interactions refer to either native gates or logical operations implemented by projective measurements. As native gates these are easy to implement and engineer in Majorana computers [2], ion traps [10], silicon spin qubits [11]. As logical operations, the interactions are implemented, for example, by lattice surgery on topological QEC (e.g. surface code). In other words, we present a very efficient way of emulating the all-to-all connectivity (on linear nearest-neighbour architectures – error-corrected or not) required for executing exponentiated Pauli strings.
@@
Figure 1: An $n$-qubit Pauli exponential is decomposed into a constant depth sequence of XX and ZZ interactions and $\mathcal{O}(n)$ ancillae. We use the shorthand notation $e^{i\theta P}=P(\theta)$, and consider that $P$ is formed exclusively of $Z$ terms: $P=Z^{\otimes n}$. A general string can always be decomposed into single-qubit Clifford gates conjugating the $Z$-terms of $P$ [9]. In this figure, we assume that $\theta$ is arbitrary, such that our construction is extending $P$ with an additional $Z$ term (marked green) and an ancilla initialized in |$\ket{\theta}=Rz(\theta)\ket{+}$ [4].
@@ I.1 Novelty
In this work: 1) we offer exact numbers (in contrast to [14, 15]) for the depth of general Pauli exponentials; 2) introduce a new set of circuit rewrite rules; 3) present the implementation of our method for fault-tolerant computers, computers with restricted topologies as well as qudit computers; 4) discuss the implications on fault-tolerant decoding, offering a novel path towards the approach presented by [16].
@@
Our decomposition is applicable to both NISQ and error-corrected machines. For the latter, we are still assuming that each logical operation is implemented with $d$ rounds of error correction, and the constant-depth refers to the computation’s space-time volume [18], which is agnostic of the code’s distance. This does not exclude the possibility of implementing Clifford logical operations in $\mathcal{1}$ rounds of error correction [20] based on improved decoding (which our method enables, but is left for future work).
A result concerning long range CNOTs has been proposed by [22]. Their adaptive circuit construction uses the same number of ancillae as our decomposition. Nevertheless, due to the exclusive focus on long-range CNOT gates, it does not solve the issue of the Clifford-ladder [14, 9]. Another constant-depth implementation of long range CNOTs using pairwise measurements is presented in [19], and is motivated by the fact that efficient decompositions of Pauli exponentials were not known at the time. Our work bridges the gap that motivated [19].
Our method is complementary to the works of [14, 19, 15]. Therein, the focus was the standard quantum circuit view of computation based on Clifford gates (e.g. CNOTs) and classical measurements, in order to implement low-depth computations using Bell pairs and more general GHZ states. However, we are not using the trick of decomposing [14, 19] or deforming circuits [15] into an entangling part followed by a computation/gate part, and a classical processing part – and approach which is similar to works such as [23, 24, 25]. In contrast, we take a different approach and treat Pauli strings as a native gate set, specific for some quantum computing architectures and general fault-tolerant computations.
@@ II Methods
The lhs. of Fig. 1 represents the unitary obtained by the exponentiation. The rhs. is the decomposition which includes ancillae for enabling the XX interactions, as well as the |$\ket{\theta}$ ancilla initialized for performing a teleportation-based gate. In all the following figures, blue boxes represent Pauli strings that include Z terms, and orange boxes are Pauli strings consisting entirely of X.
Figure 2: The ROT rewrite rule implements the exponential of the Pauli $P$ by introducing an ancilla initialized in |$\ket{\theta}=Rz(\theta)\ket{+}$, extending the Pauli string (the blue box) with a $Z$-term for the ancilla, and measuring the ancilla in the X-basis[4].
@@ II.1 Measurement-based implementation
When compiling a computation into Pauli exponentials on fault-tolerant machines, the $\theta$ rotation has to be decomposed beforehand into a sequence of rotation gates supported by the QEC [18]. This increases the computation’s depth proportionally to the length of the gate sequence approximating $\theta$. Moreover, the depth will additionally increase by a constant factor, as each rotation is probabilistic and might require corrections [4] (e.g. if the X-measurement results in the -1 eigenvalue, then a corrective rotation of $2\theta$ is needed). In the following we assume that the fault-tolerant computation has already been compiled, and when discussing the implementaiton of a logical Pauli exponential we refer to the circuit that is consuming the (approximated) rotated state |$\ket{\theta}$ (the lowest qubit in Fig. 1), and show that this circuit has a constant depth with respect to two-body XX/ZZ interactions.
Moreover, in the fault-tolerant case, we also assume that the two-body interactions are measurement-based. The latter will introduce corrective terms (e.g. [3]), but the corrections can be tracked in a Pauli frame, as the underlying circuit will consist solely of CNOT gates and ancillae initialised or measured in a discrete set of rotated basis [23, 24].
@@ II.2 Decomposing the Pauli exponentials
We start by decomposing an arbitrary Pauli string $P$ into a string $C=C_{0}\otimes\ldots\otimes C_{n}$ of single qubit Clifford gates $C_{i}$, and a string $P_{z}=Z^{\otimes n}$, such that $P=CP_{z}C^{\dagger}$. As an example, single qubit Clifford gates can be efficiently implemented in practice [26, 27] in surface codes. We continue by decomposing $P_{z}$ via the PG rule [9] (Fig. 3). CNOTs are expressed in lattice surgery style[28] decompositions (Fig. 4). After using PG, we choose the first LS rewrite for the lhs. CNOT, and the second rewrite for the rhs. CNOT.
@@
We replace a sequence of measurement and reset operations in the same basis with an equivalent measurement-based implementation by using the MR rule (Fig. 5). This is, to the best of our knowledge, a novel quantum circuit rewrite rule.
@@ III.1 Lattice Surgery
Lattice surgery is the de facto standard way of implementing computations with QLDPC codes [6], such as the surface code. Therein, the logical qubits of the circuits are arranged on a 2D layout of patches (e.g. Fig. 7), and ancillary space between the patches supports long range multi-body measurements. The latter are expressed as Pauli exponentials. The structure of the layout plays a significant role in the trade-off between a computation’s speed and the required amount of physical hardware [31, 4]. Although our decomposition uses ancillae, these are readily available in most of the lattice surgery layouts considered for compiling fault-tolerant computations.
Figure 7: Multi-body $Z^{\otimes n}$ interaction implemented by lattice surgery. The patch layout is adapted from  [31]: pink patches (Q) hold logical qubits, and gray patches (A) are ancilla for XX and ZZ interactions. The blue boundaries are the Z operators of the logical A patches. a) according to Fig. 1, we initialize twelve ancilla A in |$\ket{0}$ and perform XX two-body interactions (orange) with the Q patches; b) afterwards we perform pairwise ZZ two-body interactions (blue) between pairs of Q and A patches. Finally, the interactions from a) are repeated. The interaction schedule from a) and b) is highly regular and can be further optimized in terms of hardware overheads.
The general compilation strategy for lattice surgery is to first decompose arbitrary algorithms into Pauli exponentials [4], and then to implement these exponentials via multi-body measurements. This approach has multiple disadvantages.
@@ III.2 Parallel Decoding
Our decomposition enables the parallelisation of decoding by setting an upper bound on the maximum (equivalent) decodable distance that a QEC decoder will need to handle within one QEC cycle. In the following, we introduce the concept of equivalent decodable distance.
@@
At the same time, as seen in Fig. 7, all blue patches can be decoded in parallel. Our parallelisation method is reminiscent of the proposal from [34] (fixing the number of physical qubits per decoder) and the constructions from [35, 36] (decoders which communicate messages). For the latter, the decoders will communicate the observed correlated logical errors [20] after the long-range interaction is split via lattice surgery. Our method offers the same decoding benefits like [16], with the caveat that we generate automatically a regular pattern of logical blocks, our the size of our blocks is upper bounded.
@@ III.3 Constant Depth Fault-Tolerant Protocols
From the perspective of stabilizer QECC codes, our decomposition allows any set of QECC stabilizers to be implemented in constant depth using pairwise XX and ZZ measurements. This claim follows for Z-strings from Fig. 1 and 3) and requires two additional depth-one layers of single qubit Clifford gates (for a total depth-five construction) to achieve arbitrary stabilizers.
QECC code conversion schemes [37] can also be implemented locally using pairwise interactions. Additionally, after carefully choosing the pairwise interaction schedules, our method might lead to results similar to the ones from [38]. Our decomposition involves only two-body measurements and may find relevance to subsystem codes such as the Bacon-Shor code [39] or its Floquet version [38].
@@ IV Generalization
So far, we considered constant depth decompositions of Pauli strings of the form $Z^{\otimes n}$. Nevertheless, there might be qubits which are not involved in the operation, and the corresponding Pauli string will include an $I$ term for those qubits.
@@ IV.1 Qudits
If ququarts are not an option, qubits can be used to simulate the effect of ququarts. Using Fig. 10a, one can show that $e^{i\theta P_{1}}\ket{A\ldots B\ldots C}=e^{i\theta P_{2}}\ket{A\ldots 0\ldots C}$ for $P_{1}=Z\ldots I\ldots Z$ and $P_{2}=Z\ldots Z\ldots Z$. This means that, whenever the Pauli string includes an $I$ on a particular qubit (e.g. $B$) then it is possible to swap $B$ with a |$\ket{0}$ and to have a Pauli string that includes a $Z$ term instead of $I$. Therefore, an alternative to using ququarts is to use one additional qubit $q$ placed next to each ancilla $a$ from Fig. 1. These $q$ qubits will be initialized in |$\ket{0}$, and the $I$ operators from the Pauli string would be replaced by $SWAP(a,q)$, a $Z$ on $a$, followed by $SWAP(a,q)$ (Fig. 8). This construction would not be nearest-neighbour, but next-nearest neighbour.
Figure 8: A next-nearest neighbour setup in order to implement the Pauli string in constant depth: using an ancilla initialized in |$\ket{0}$ and swapped before and after the Pauli string.
@@ IV.2 The qubit layout
In a linear nearest neighbour layout of qubits, the logical qubits are arranged in a line similar to $\begin{smallmatrix}Q&0&Q&0\ldots&Q\end{smallmatrix}$, where 00 are ancillae initialized in |$\ket{0}$. In this scenario, pairwise nearest-neighbour interactions are not possible without resorting to ququarts, as discussed in the previous section. Consequently, the long range construction from [22] uses nearest-neighbour pairwise interactions, but it is limiting the state of the intermediate qubits to the less general layout $\begin{smallmatrix}Q&0&0&0\ldots&Q\end{smallmatrix}$ – all the qubits between the first and last one are initialized in |$\ket{0}$.
Yet another solution to achieving constant depth implementation of the Pauli string is to use a bi-linear nearest neighbour layout of qubits of the form $\begin{smallmatrix}Q&0&Q&0\ldots&Q\\ r&r&r&r\ldots&r\end{smallmatrix}$. This layout allows for jumping over qubits operated by $I$ and maintaining the nearest-neighbour decomposition. If the $r$ qubits would be operated in a fault-tolerant manner, protected by surface codes, these can be used to first prepare Bell pairs using the constant depth recipe from Fig. 18 in [19] and then to follow the Pauli string implementation using Bell pairs [19, 14, 15]. Simultaneously, the layout from Fig. 7 has the same properties the bi-linear nearest neighbour layout has.